# Citadel: Context Similarity Based Deep Learning Framework Bug Finding

### v4

@@ 5.2. Effectiveness in Detecting Bugs
Rejected Case: Although Citadel identifies new bugs based on whether analogous APIs exhibit the same anomalous behavior as the problematic API, several cases are still rejected by developers. Note that, these rejected cases still exhibit anomalous behaviors, but developers consider them unimportant and have no plans to fix or work on them. Here, we present an example. Citadel encounters a rejected case when detecting a bug on the torch.logdet API. During testing, Citadel matches torch.logdet with torch.det and then constructs test cases for the target API logdet based on a buggy code of the source API det. When executing the test cases, Citadel finds that both APIs have abnormal and dangerous output value ‘NaN’ (i.e., not a number) which could further affect subsequent calculations and raise dangerous behaviors (Odena et al., 2019). However, developers suggest that the abnormal ‘NaN’ output on logdet is due to the calculation characteristics of this API. When the input matrix is close to being non-invertible (e.g., very small singular values), then these APIs may return incorrect results (tor, 2023a). Rejected cases show the limitations of Citadel in identifying bugs on APIs that are allowed to exhibit anomalous behaviors. How to identify and filter these unimportant anomalous behaviors to reduce rejected cases will be our future direction.
Answer to RQ1: Citadel can effectively utilize the bug reports of source APIs to identify new bugs in their analogous API functions. In the experiment, Citadel leverages 172 bug cases to detect 151 API bugs, including 13 performance bugs that existing methods fail to capture, demonstrating the effectiveness of Citadel in finding various real wolrd bugs. Furthermore, 49 of these detected bugs originate from analogous API pairs that existing approaches cannot match, while 9 bugs are from APIs that existing methods do not cover. These results highlight the contribution of the APIs and API pairs newly covered by Citadel in enhancing bug detection.
@@ 5.3. Efficiency in Detecting Bugs
Experiment Design and Results: To evaluate the efficiency of Citadel in generating test cases and triggering bugs, we conduct experiments to calculate and compare the Ratio of test cases that can trigger bugs and Average time to detect bugs of Citadel and baseline methods. Specifically, we execute the complete procedure of Citadel and three baselines to record the total number of generated test cases and the number of test cases that can trigger bugs, and the time cost of testing, as described in § 5.1. As the baselines may not save some test cases or inputs, to ensure that we do not mistakenly count such excluded test cases, we directly record the total number of generated test files as the total amount of test cases. The experimental results are shown in Table 2. The column ‘#Valid’ shows the number of generated cases that can trigger bugs and the column ‘#Total’ denotes the overall number of generated cases in one complete execution. The column ‘Ratio’ displays the ratio of the number of cases that triggered bugs to the total number of generated test cases, corresponding to the ‘Ratio of test cases that can trigger bugs’ in § 5.1. The last column of Table 2 displays the ‘Average time to detect bugs’ of each approach.
@@
Analysis: The results in Table 2 demonstrate the efficiency of Citadel in generating test cases to trigger bugs. In the complete execution of baselines, DocTer generates 33,859 test files on PyTorch and TensorFlow in a complete execution, and only 251 of them (0.74% of the total) are valid cases. The testing process on two frameworks lasts over 99 hours, and the average time to detect bugs in DocTer is 41.98 min. On the PyTorch framework, DeepREL spends over 27 hours generating a total of 77,662 test files and marks 2,001 of them as ‘can trigger bugs’, and the valid test case ratio is 2.58%. DeepREL does not provide the test cases that can trigger bugs or the corresponding number on TensorFlow. Therefore, we count the number of candidate bugs it detects (the upper bound of possible bug cases) in Table 2. DeepREL spends over 150 hours testing two frameworks, and at most 1.23% of all 330,195 generated test files can trigger bugs. The average time to detect bugs of DeepREL is 58.62 min. TitanFuzz spends over 72 hours generating 350,047 cases on two DL frameworks, 3.90% of which are candidates and can catch valuable buggy behaviors or trigger bugs. The average time of TitanFuzz to detect bugs is 68.43 minutes. By contrast, Citadel average spends 5.70 min to detect one bug, which is only 13.57%, 9.72%, and 8.33% of the average time cost of DocTer, DeepREL, and TitanFuzz, respectively. Citadel is over 10x more time efficient than DeepREL and TitanFuzz in detecting bugs. In testing, Citadel generates a total of 404 test files based on 172 collected bugs within 15 hours, and 143 of them can used to discover bugs and the ratio reaches 35.40%. Note that quite a part of generated test files can be used to detect bugs for multiple analogous API functions at the same time, which improves the efficiency of Citadel in generating test cases and detecting bugs. In addition, our experiments show that the time from when Citadel starts testing to when it triggers the first bug is within 3 minutes, while baselines usually tend to take 8 minutes or more. Citadel, therefore, outranks three baselines in generating and utilizing test cases to detect bugs efficiently.
@@
As a bug-finding tool that is orthogonal to existing tools, Citadel can also be used to enhance the effectiveness and efficiency of bug detection. In the experiment, Citadel leverages 13 bug cases from the baseline method reports and successfully detects 18 API bugs within 3 hours, 11 of which are not reported by the baselines. This demonstrates the potential of collaboration between Citadel and existing bug-finding tools to accelerate DL framework bug detection.
@@
Answer to RQ2: Citadel can efficiently generate test cases to trigger new bugs. In the experiments, 35.40% of the test cases generated by Citadel can trigger bugs, which is over 9.08 times the bug triggering ratio of the baseline methods. Furthermore, the average time to detect bugs of Citadel is 5.7 minutes, while baseline methods need at least 41.98 minutes, demonstrating the efficiency of Citadel in detecting real-world bugs.
@@ 6. Related Work
DL Framework Testing. Researchers have proposed various methods to test DL framework (e.g., PyTorch and TensorFlow) through model-level methods (Pham et al., 2019; Guo et al., 2020; Wang et al., 2020b; Gu et al., 2022) and API-level methods (Wang et al., 2022; Deng et al., 2022; Wei et al., 2022; Xie et al., 2022), which have been comprehensively introduced in § 2.2. In addition, existing works also design elaborated metamorphosis relations to validate the correctness of DL framework implementation (Ding et al., 2017; Wang et al., 2020a). Recently, Zhang et al. (Zhang et al., 2021) propose the test tool ‘Predoo’, which performs a fine-grained evaluation of the shape variable input and error of 7 operators of the DL framework. Researchers also focus on the security problem of DL frameworks. SkipFuzz (Kang et al., 2022) uses active learning to learn the input constraints of different library API functions and generates valid test inputs for TensorFlow and PyTorch. It had finally identified 43 crashes on DL frameworks, including 13 CVEs assigned. IvySyn (Christou et al., 2023) constructs code blocks by DL framework APIs based on a set of offending inputs that trigger memory safety errors in the underlying implementation of DL frameworks (e.g., in C/C++ program language) to trigger security vulnerabilities. In addition, researchers also detect bugs in other DL underlying libraries such as DL compilers. Liu et al. (Liu et al., 2022) design testing method for ML compiler framework TVM. This method is guided by coverage feedback to mutate the low-level intermediate representation of TVM to achieve more effective fuzzing testing. Shen et al. (Shen et al., 2024) transfer the knowledge of DL framework fuzzers (e.g., DocTer and DeepRel) to generate effective test cases for diverse DL compiler operators and detect crashes and inconsistencies. Recently, Zhang et al. (Zhang et al., 2025) survey the testing methods on various DL libraries and point out the challenges of future testing research.
@@
Different from the above testing methods, Citadel is built on the concept of code similarity measurement and real-world bugs confirmed by developers. Citadel focuses on DL framework bugs and can leverage existing bug reports on one API to efficiently exploit new bugs on its analogous APIs.
@@
Code Similarity Measurement. Existing work has a variety of code similarity measurement and code clone detection methods. These methods can be divided into static and dynamic according to whether the test code needs to be executed. The static methods mainly include metrics-based (Ottenstein, 1976; Faidhi and Robinson, 1987), text-based (Roy and Cordy, 2008; Luo et al., 2014), token-based (Sajnani et al., 2016; Đurić and Gašević, 2013), AST-based (Zhang et al., 2019; Jiang et al., 2007), and graph-level methods (Crussell et al., 2012; Chae et al., 2013). Among them, the AST and graph-level methods can comprehensively understand the syntax and capture the relationship of the function calls between code blocks, therefore, they have better detection effects but greater overhead than other methods. Existing dynamic methods propose that the functional similarity between programs can be evaluated from the perspectives of input and output (Elva and Leavens, 2012; Su et al., 2016), abstract memory state (Kim et al., 2011), etc. Recently, Maertens et al. (Maertens et al., 2022) propose an open-source tool that supports a broad range of programming languages (e.g., C, Java, Python, go). Zhang et al. (Zhang et al., 2022b) present an AST-assisted approach for generalizable neural clone detection to find clones in codebases reflecting industry practices. Wang et al. (Wang et al., 2023) design a code clone detection tool based on the semantic token which enhances the detection capability by complementing the traditional token with semantic information. Our work mainly combines the concept of code similarity measurement into DL framework testing. Citadel matches context-similar APIs based on the similarity of their context information and leverages the bug reports on one API to efficiently find new bugs on its analogous APIs.
@@ 7. Discussion
Citadel Enhancement.Automated Verification and Annotation: Currently, Citadel receives the bug case list as one of its inputs, which requires manual intervention. How to automatically verify and annotate bug cases (e.g., using code models) is a potential future direction to enhance Citadel’s efficiency. ❷ Performance Bug Reports: Citadel presently handles performance bug reports that provide code snippets to estimate expected overhead. A valuable future direction would involve automating the calculation of expected overhead or extracting expected behavior from bug reports described in natural language or images (tor, 2022c). ❸ Fault Localization: Some API bugs may have the same root cause, and one effective patch may fix multiple API bugs (tor, 2022a). The detection results and fix patches of prior work (tor, 2020) also support this observation. Considering that Citadel, like existing bug-finding tools (Xie et al., 2022; Deng et al., 2022), is designed for bug detection rather than faulty localization and program repair, determining the root cause of API bugs is out of scope. We have reported the number of duplicated bugs in Table 1. Designing bug localization methods to localize and fix the detected bugs will be a future direction. ❹ Code Similarity Measurement: To mitigate false positives in the static analyzer, Citadel selects strict thresholds to judge analogous source code functions based on the experimental results of prior work (Su et al., 2016; Misu and Sakib, 2017). Even if there are still a few false positives in the static analyzer, they can hardly affect the subsequent API matching and bug detection. On the one hand, API functions that are mistakenly matched due to these false positives are easily discarded by the API matcher due to argument mismatches. On the other hand, Citadel identifies bugs when the target API exhibits the same buggy behavior that is consistent with the source API’s bug report. Therefore, Citadel can always detect real anomalous behavior. However, the static analyzer could still miss some analogous source code functions. Specifically, different developers may use the same logic to construct functions to solve specific problems, resulting in different implementations of two source code functions but similar functionality (Zakeri-Nasrabadi et al., 2023). Currently, the static analyzer cannot effectively identify such analogous functions. How to enhance the matching effect of similar source code functions and APIs in Citadel will be a future direction.
@@
DL Bug Finding.LLM Library Bug Finding: With the advent of Large Language Model (LLM) technology, dependent libraries (e.g., APEX (ape, 2024)) of LLMs exhibit various bugs, such as crashes during model training (Zhang et al., 2022a), which impede the application and deployment of LLMs. However, the substantial runtime overhead of LLMs renders traditional DL fuzz testing methods, which generate millions of test cases, unsuitable for LLM library testing. Developing an efficient LLM library testing method to uncover potential bugs will be a valuable future direction. ❷ Performance Bug Detection: Current DL framework testing primarily addresses crashes and numerical inconsistencies, with little attention given to performance bugs that impact model training, economy, and even the environment. Detecting performance bugs necessitates constructing a test oracle and estimating the expected overhead. Our finding suggests that different settings of certain API arguments may influence the actual overhead (e.g., ‘Bug Case 1’ of § 5.2). Therefore, the runtime overhead of one API function can be qualitatively or even quantitatively estimated based on the description of these arguments, serving as a pseudo-test oracle. Formalizing expected performance changes from API argument descriptions and constructing test oracles to detect performance bugs will be an important future direction.

### v5

@@ 5.2. Effectiveness in Detecting Bugs
Rejected Case: Although Citadel effectively identifies API bugs based on whether analogous APIs exhibit the same anomalous behavior as the problematic API, several cases are still rejected by developers. Note that, these rejected cases still exhibit anomalous behaviors, but developers consider them unimportant and have no plans to fix or work on them. Here, we present an example. Citadel encounters a rejected case when detecting a bug on the torch.logdet API. During testing, Citadel matches torch.logdet with torch.det and then constructs test cases for the target API logdet based on a buggy code of the source API det. When executing the test cases, Citadel finds that both APIs have abnormal and dangerous output value ‘NaN’ (i.e., not a number) which could further affect subsequent calculations and raise dangerous behaviors (Odena et al., 2019). However, developers suggest that the abnormal ‘NaN’ output on logdet is due to the calculation characteristics of this API. When the input matrix is close to being non-invertible (e.g., very small singular values), then these APIs may return incorrect results (tor, 2023a). Rejected cases show the limitations of Citadel in identifying bugs on APIs that are allowed to exhibit anomalous behaviors. How to identify and filter these unimportant anomalous behaviors to reduce rejected cases will be our future direction.
Answer to RQ1: Citadel can effectively utilize the bug reports of source APIs to identify API bugs in their analogous API functions. In the experiment, Citadel leverages 172 bug cases to detect 151 API bugs, among which 124 are newly reported and 13 are performance bugs that existing methods fail to detect. These results demonstrate the effectiveness of Citadel in finding a wide range of real-world bugs. Furthermore, 49 of these detected bugs originate from analogous API pairs that existing approaches cannot match, while 9 bugs are from APIs that existing methods do not cover. These results highlight the contribution of the APIs and API pairs newly covered by Citadel in enhancing bug detection.
@@ 5.3. Efficiency in Detecting Bugs
Experiment Design and Results: To evaluate the efficiency of Citadel in generating test cases and triggering bugs, we conduct experiments to calculate and compare the Ratio of test cases that can trigger bugs and Average time to detect bugs of Citadel and baseline methods. Specifically, we execute the complete procedure of Citadel and three baselines to record the total number of generated test cases and the number of test cases that can trigger bugs, and the time cost of testing, as described in § 5.1. As the baselines may not save some test cases or inputs, to ensure that we do not mistakenly count such excluded test cases, we directly record the total number of generated test files as the total number of test cases. The experimental results are shown in Table 2. The column ‘#Valid’ shows the number of generated cases that can trigger bugs and the column ‘#Total’ denotes the overall number of generated cases in one complete execution. The column ‘Ratio’ displays the ratio of the number of cases that triggered bugs to the total number of generated test cases, corresponding to the ‘Ratio of test cases that can trigger bugs’ in § 5.1. The last column of Table 2 displays the ‘Average time to detect bugs’ of each approach.
@@
Analysis: The results in Table 2 demonstrate the efficiency of Citadel in generating test cases to trigger bugs. In the complete execution of baselines, DocTer generates 33,859 test files on PyTorch and TensorFlow in a complete execution, and only 251 of them (0.74% of the total) are valid cases. The testing process on two frameworks lasts over 99 hours, and the average time to detect bugs in DocTer is 41.98 minutes. On the PyTorch framework, DeepREL spends over 27 hours generating a total of 77,662 test files and marks 2,001 of them as ‘can trigger bugs’, and the valid test case ratio is 2.58%. DeepREL does not provide the test cases that can trigger bugs or the corresponding numbers on TensorFlow. Therefore, we count the number of candidate bugs it detects (the upper bound of possible bug cases) in Table 2. DeepREL spends over 150 hours testing two frameworks, and at most 1.23% of all 330,195 generated test files can trigger bugs. The average time to detect bugs of DeepREL is 58.62 minutes. TitanFuzz spends over 72 hours generating 350,047 cases on two DL frameworks, 3.90% of which are candidates and can catch valuable buggy behaviors or trigger bugs. The average time of TitanFuzz to detect bugs is 68.43 minutes. By contrast, Citadel average spends 5.70 minutes to detect one bug, which is only 13.57%, 9.72%, and 8.33% of the average time cost of DocTer, DeepREL, and TitanFuzz, respectively. Citadel is over 10x more time efficient than DeepREL and TitanFuzz in detecting bugs. In testing, Citadel generates a total of 404 test files based on 172 collected bugs within 15 hours, and 143 of them can be used to discover bugs, and the ratio reaches 35.40%. Note that quite a part of the generated test files can be used to detect bugs for multiple analogous API functions at the same time, which improves the efficiency of Citadel in generating test cases and detecting bugs. In addition, our experiments show that the time from when Citadel starts testing to when it triggers the first bug is within 3 minutes, while baselines usually tend to take 8 minutes or more. Citadel, therefore, outranks three baselines in generating and utilizing test cases to detect bugs efficiently.
@@
Existing bug-finding tools use various techniques, such as fuzzing, to explore and discover new anomalous behaviors and detect bugs. As a bug-finding tool orthogonal to existing tools, Citadel aims to use reported bugs to discover API bugs on analogous APIs, therefore, it can be used to enhance the effectiveness and efficiency of bug detection. In the experiment, Citadel leverages 13 bug cases from the baseline method reports and successfully detects 18 API bugs within 3 hours, 11 of which are not reported by the baselines. This demonstrates the potential of collaboration between Citadel and existing bug-finding tools to accelerate DL framework bug detection.
@@
Answer to RQ2: Citadel can efficiently generate test cases to trigger API bugs. In the experiments, 35.40% of the test cases generated by Citadel can trigger bugs, which is over 9.08 times the bug triggering ratio of the baseline methods. Furthermore, the average time to detect bugs of Citadel is 5.7 minutes, while baseline methods need at least 41.98 minutes, demonstrating the efficiency of Citadel in detecting real-world bugs.
@@ 6. Related Work
DL Framework Testing. Researchers have proposed various methods to test DL framework (e.g., PyTorch and TensorFlow) through model-level methods (Gu et al., 2022; Mu et al., 2025b; Wang et al., 2024) and API-level methods (Wang et al., 2022; Deng et al., 2022; Xie et al., 2022), which have been comprehensively introduced in § 2.2. In addition, existing works also design elaborated metamorphic relations to validate the correctness of DL framework implementation (Ding et al., 2017; Wang et al., 2020a). FreeFuzz (Wei et al., 2022) successfully detects one performance bug using metamorphic testing techniques. However, limited by its metamorphic relation, it can only test framework behaviors related to tensor types and cannot effectively identify other diverse performance bugs (e.g., the LazyConvTranspose2d bug in Fig. 2 and the Hardtanh bug in Fig. 7). Recently, Zhang et al. (Zhang et al., 2021) propose the test tool ‘Predoo’, which performs a fine-grained evaluation of the shape variable input and error of 7 operators of the DL framework. Researchers also focus on the security problem of DL frameworks. SkipFuzz (Kang et al., 2022) uses active learning to learn the input constraints of different library API functions and generates valid test inputs for TensorFlow and PyTorch. It had finally identified 43 crashes on DL frameworks, including 13 CVEs assigned. IvySyn (Christou et al., 2023) constructs code blocks by DL framework APIs based on a set of offending inputs that trigger memory safety errors in the underlying implementation of DL frameworks (e.g., in C/C++ program language) to trigger security vulnerabilities. In addition, researchers also detect bugs in other DL underlying libraries such as DL compilers. Liu et al. (Liu et al., 2022) design a testing method for the ML compiler framework TVM. This method is guided by coverage feedback to mutate the low-level intermediate representation of TVM to achieve more effective fuzzing testing. Shen et al. (Shen et al., 2024) transfer the knowledge of DL framework fuzzers (e.g., DocTer and DeepRel) to generate effective test cases for diverse DL compiler operators and detect crashes and inconsistencies. Shiri et al. (Shiri Harzevili et al., 2025) extract corner case patterns from historical issue reports to guide fuzzers in generating test cases and finding bugs more effectively. Their method uses differential testing and crashes to construct test oracles and effectively identifies status and value bugs. However, this design still suffers from the limitation described in § 3, namely, the lack of capability of detecting performance bugs. To effectively test rapidly iterating DL frameworks, Xie et al. (Xie et al., 2024) built upon previous work (Xie et al., 2022; Wang et al., 2022) to design a continuous testing framework for efficiently discovering regression bugs and masked bugs. Recently, Zhang et al. (Zhang et al., 2025) survey the testing methods on various DL libraries and point out the challenges of future testing research.
@@
Different from the above methods, Citadel is built on the concept of code similarity measurement and real-world bugs confirmed by developers. Citadel focuses on DL framework bugs and can leverage existing bug reports on one API to efficiently exploit bugs on its analogous APIs.
@@
Code Similarity Measurement. Existing work has a variety of code similarity measurement and code clone detection methods. These methods can be divided into static and dynamic according to whether the test code needs to be executed. The static methods mainly include metrics-based (Ottenstein, 1976; Faidhi and Robinson, 1987), text-based (Roy and Cordy, 2008; Luo et al., 2014), token-based (Sajnani et al., 2016; Đurić and Gašević, 2013), AST-based (Zhang et al., 2019; Jiang et al., 2007), and graph-level methods (Crussell et al., 2012; Chae et al., 2013). Among them, the AST and graph-level methods can comprehensively understand the syntax and capture the relationship of the function calls between code blocks, therefore, they have better detection effects but greater overhead than other methods. Existing dynamic methods propose that the functional similarity between programs can be evaluated from the perspectives of input and output (Elva and Leavens, 2012; Su et al., 2016), abstract memory state (Kim et al., 2011), etc. Recently, Maertens et al. (Maertens et al., 2022) propose an open-source tool that supports a broad range of programming languages (e.g., C, Java, Python, go). Zhang et al. (Zhang et al., 2022b) present an AST-assisted approach for generalizable neural clone detection to find clones in codebases reflecting industry practices. Wang et al. (Wang et al., 2023) design a code clone detection tool based on the semantic token which enhances the detection capability by complementing the traditional token with semantic information. Our work mainly combines the concept of code similarity measurement into DL framework testing. Citadel matches context-similar APIs based on the similarity of their context information and leverages the bug reports on one API to efficiently find new API bugs on its analogous APIs.
@@ 7. Discussion
Discussion.Impact of Collected Bug Cases: Different from existing fuzzing methods (Xie et al., 2022; Deng et al., 2023), which directly generate or mutate test cases for individual APIs, Citadel generates new test cases and finds API bugs for analogous APIs based on collected bug cases. Note that the detection results of Citadel are inherently dependent on the collected bug cases. The more extensive and diverse problematic APIs these cases encompass, the more comprehensive the API coverage Citadel can achieve during detection, leading to the discovery of additional previously unknown bugs. Conversely, a limited number of bug cases concentrated on specific APIs constrain Citadel to testing only a narrow subset of analogous APIs. Although the experiments in § 5.2 successfully cover 1,204 DL framework APIs and detect 124 unreported API bugs based on the 172 collected bug cases, demonstrating the effectiveness and efficiency of Citadel, this still represents only a small fraction of the matched analogous APIs, as shown in Table 2. Developing techniques to automatically collect and annotate a more diverse set of bug cases, thereby finding bugs on a broader range of APIs, is a promising future research direction. ❷ Scope of Citadel: As a testing tool orthogonal to existing bug-finding approaches, rather than searching for unknown API bugs from scratch, Citadel leverages a known bug in one API to uncover bugs in its similar APIs. Note that the API bugs discovered by Citadel may stem from the same underlying implementation errors as the known API bugs. However, this does not diminish the value of Citadel. Effectively identifying potential API bugs is crucial for the development and maintenance of DL frameworks. For framework users, Citadel could promptly recognize APIs exhibiting abnormal behavior, thereby preventing users from misusing these buggy APIs and mitigating potential security risks in users’ models and software. For framework developers, Citadel efficiently reveals abnormal behaviors across a series of APIs that share similar underlying implementations, providing developers a more comprehensive debugging perspective, helping them infer the root cause and efficiently localize and fix bugs. Moreover, the experiment results in § 5.3 demonstrate that Citadel utilizes 13 bug cases reported by baselines to detect 18 API bugs within three hours, 11 of which are previously unreported, demonstrating the efficiency and practicality of Citadel. These findings highlight that Citadel is not a replacement for existing methods, but rather a complement to existing testing tools. It can effectively leverage individual bug reports and improve the effectiveness and efficiency of the entire bug detection and repair process.
@@
Citadel Enhancement.Automated Verification and Annotation: Currently, Citadel receives the bug case list as one of its inputs, which requires manual intervention. How to automatically verify and annotate bug cases (e.g., using code models) is a potential future direction to enhance Citadel’s efficiency. ❷ Performance Bug Reports: Citadel presently handles performance bug reports that provide code snippets to estimate expected overhead. A valuable future direction would involve automating the calculation of expected overhead or extracting expected behavior from bug reports described in natural language or images (tor, 2022c). ❸ Fault Localization: Some API bugs may share a common root cause, and a single effective patch can potentially resolve multiple related issues (tor, 2022a). The detection results and corresponding fix patches reported in prior work (Xie et al., 2022; tor, 2020) further support this observation. Note that Citadel, like existing bug-finding tools (Xie et al., 2022; Deng et al., 2022), is designed for bug detection rather than faulty localization or program repair. As such, identifying the root causes of API bugs lies outside the scope of this testing work. We have reported the number of duplicate bugs and the number of API bugs that share the same implementation errors as the source bugs in § 5.2. However, since our reports for TensorFlow API bugs currently do not receive any patches, we are unable to conduct the analysis on this framework. We estimate that API bugs with shared underlying implementations also exist in this framework. Therefore, designing error localization methods for automated fault localization remains a valuable direction for future research, which could effectively reduce the debugging workload for DL framework developers (e.g., TensorFlow) and improve bug repair efficiency. ❹ Code Similarity Measurement: To mitigate false positives in the static analyzer, Citadel selects strict thresholds to judge analogous source code functions based on the experimental results of prior work (Su et al., 2016; Misu and Sakib, 2017). Even if there are still a few false positives in the static analyzer, they can hardly affect the subsequent API matching and bug detection. On the one hand, API functions that are mistakenly matched due to these false positives are easily discarded by the API matcher due to argument mismatches. On the other hand, Citadel identifies bugs when the target API exhibits the same buggy behavior that is consistent with the source API’s bug report. Therefore, Citadel can always detect real anomalous behavior. However, the static analyzer could still miss some analogous source code functions. Specifically, different developers may use the same logic to construct functions to solve specific problems, resulting in different implementations of two source code functions but similar functionality (Zakeri-Nasrabadi et al., 2023). Currently, the static analyzer cannot effectively identify such analogous functions. How to enhance the matching effect of similar source code functions and APIs in Citadel will be a future direction. ❺ Integration with Existing Tools: The detection effectiveness of Citadel is influenced by the diversity and quality of the collected bug cases. Therefore, as described in § 4.1, we collect and validate bug cases in the preparation stage to ensure the quality of the bug cases extracted from GitHub repositories. A promising future direction is to integrate Citadel with other bug-finding tools (e.g., fuzzing) into a unified testing pipeline. In such a pipeline, once an API bug is detected by an external tool, the problematic API and its bug-triggering test case can be directly used as input to Citadel. Citadel can then efficiently construct new test cases and detect bugs in analogous APIs. This workflow has the potential to substantially reduce manual effort, enhance the efficiency of bug detection, and assist developers in more effectively identifying the root causes of bugs.
@@
DL Bug Finding.LLM Library Bug Finding: With the advent of Large Language Model (LLM) technology, dependent libraries (e.g., APEX (ape, 2024)) of LLMs exhibit various bugs, such as crashes during model training (Zhang et al., 2022a; Mu et al., 2025a; Jiang et al., 2025), which impede the application and deployment of LLMs. However, the substantial runtime overhead of LLMs renders traditional DL fuzz testing methods, which generate millions of test cases, unsuitable for LLM library testing. Developing an efficient LLM library testing method to uncover potential bugs will be a valuable future direction. ❷ Performance Bug Detection: Current DL framework testing primarily addresses crashes and numerical inconsistencies, with little attention given to performance bugs that impact model training, economy, and even the environment. Detecting performance bugs necessitates constructing a test oracle and estimating the expected overhead. Our finding suggests that different settings of API arguments may influence the actual overhead (e.g., ‘Bug Case 1’ of § 5.2). Therefore, the runtime overhead of one API function can be qualitatively or even quantitatively estimated based on the description of these arguments, serving as a pseudo-test oracle. Formalizing expected performance changes from API argument descriptions and constructing test oracles to detect performance bugs will be an important future direction.